### v1

@@ 6 Conclusions
The numerical experiments show that the resulting reconstruction errors are comparable to those obtained using the full-order model alone, yet at a significantly lower cost. Furthermore, since the reconstruction error does not deviate from the reference behavior as the state uncertainty decreases, we can conclude that the adaptive tolerance proposed is effective in ensuring surrogate models with the necessary accuracy.
Regarding future developments, integrating adaptive hyper-reduction techniques into the on-the-fly training phase of the surrogate model is expected to enhance the algorithm’s performance and handle non-linearity more efficiently. This, together with the implementation of the method using more than two levels of accuracy, constitutes the natural progression of the proposed approaches.

### v2

@@ 6 Conclusions
The numerical experiments show that the adaptive multi-level algorithm achieves errors comparable to those obtained using the full order model alone, yet at a significantly lower cost. Furthermore, since the reconstruction error does not deviate from the reference behavior as the state uncertainty decreases, we can conclude that the proposed adaptive tolerance is effective in ensuring that surrogate models maintain the necessary accuracy. Between the aRB-ML-EnKF and the aRB-MF-EnKF, the former appears to be a more reliable option. Indeed, although in most of the experiments the adaptive multi-fidelity approach yields similar results, the last experiment involving larger ensembles seems to suggest convergence to a different, and in this case less accurate, mean field limit.
Regarding future developments, integrating adaptive hyper-reduction techniques into the on-the-fly training phase of the surrogate model, as in donoghue22 , is expected to enhance the algorithm’s performance and handle non-linearity more efficiently. This, together with the implementation of the method using more than two levels of accuracy, constitutes the natural progression of the proposed approaches.